OpenStudy (anonymous):

Find the real values of x such that 3^(2x^2-7x+3)= 4^(x^2-x-6).

OpenStudy (anonymous):

Tricky . . . going to need a couple steps with logarithms. Factoring the exponents might help later on, so \[3^{(2x-1)(x-3)}=4^{(x-3)(x+2)}\]

OpenStudy (anonymous):

To make things a little easier to type, I'm going to make the following substitutions: a=2x-1 b=x-3 c=x+2 so 3^(ab)=4^(bc). Now take the log, base-3 of both sides.

OpenStudy (anonymous):

Can you figure out the next couple steps from there?

OpenStudy (anonymous):

how do you take the log of a different base other than 10? because it's been a while since i've worked on logs without a calcultor

OpenStudy (anonymous):

Ah, well, unless you want to calculate a decimal approximation, you can leave it in log form.

OpenStudy (anonymous):

But in general, let's say you wanted to find the log, base-3 of 4 (which shows up in this problem), you find the log, base-10 of 4 then divide by the log, base-10 of 3. That's the change-of-base formula.

OpenStudy (anonymous):

The next step would look like this: \[\log_3[3^{ab}=4^{bc}] \rightarrow ab=bc \log_3(4).\]

OpenStudy (anonymous):

The b's cancel from both sides and you can algebra your way through the rest to solve for x.

OpenStudy (anonymous):

I'll chew on this for a while and get back to you. Logs are not my strong point.

OpenStudy (anonymous):

Won't necessarily need this fact for this problem, but remember that logarithms are exponents, so follow all the same rules for exponents. Good luck! (The approximate value for x I got was around 4.77)

OpenStudy (anonymous):

Wouldn't it be simpler to do log base 10 from both sides to bring the equations down so it will end up looking like: (ab)(log3) = (bc)(log4) with log base 10? Or is that an illegal move with logs?